OpenStudy (anonymous):

How do i decrease the dimensions of a rectangular prism by 1/4?

OpenStudy (jdoe0001):

\(\bf \textit{dimension of a cube}=length\cdot width\cdot height\\ \quad \\ \textit{dimension of a Regular cube}=length\cdot width\cdot height=x\cdot x\cdot x\\ \quad \\ \frac{1}{4}\textit{ of each side?}\implies \cfrac{length}{4}\cdot \cfrac{width}{4}\cdot \cfrac{height}{4}=\cfrac{x}{4}\cdot \cfrac{x}{4}\cdot \cfrac{x}{4}\implies \cfrac{x^3}{4^3}\)

OpenStudy (anonymous):

thanx! i appreciate the help.